OpenStudy (anonymous):

Solve 2x2 = 72. A.{±18} B.{±36} C.{±6} D.{±8}

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

@jagr2713

OpenStudy (anonymous):

@paki

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

@jagr2713

Nnesha (nnesha):

first solve for x^2 move 2 to the right side

OpenStudy (anonymous):

ok

Nnesha (nnesha):

let me know what you get after that

OpenStudy (anonymous):

ok hld on

OpenStudy (anonymous):

ok so are we moving 2 this way 72*2x?

Nnesha (nnesha):

2 is multiplying with x^2 so to cancel 2 from left side you should divide 2 both sides

OpenStudy (anonymous):

1*36

OpenStudy (anonymous):

so its B

Nnesha (nnesha):

how ?

Nnesha (nnesha):

x^2 =36 not x

OpenStudy (anonymous):

well 2/72=36 then 2/2=1*36=36

Nnesha (nnesha):

yes right so x^2 =36 now to cancel out 2 take square root both sides

OpenStudy (anonymous):

that would be 72 square root 36 by 2

Nnesha (nnesha):

nope  72/2 is 36 so then there isn't any 72 to take square root of 72